(4 pts) A histogram of monthly wages for part-time employees is shown below. No one earned more than 700 dollars a month. AnsPlease help me solve this and also show all working. Thanks.
Answer #1

In above Histogram, x-axis represents the Wages (hundred of dollars) For ex. 1 mean 100 dollars, 2 means 200 dollars etc.

And y-axis represents percentage (%) per $100. For ex. 25 means 25%.

Using this information we get percentage for the following cases:

Let X be a wages (hundred of dollars)

i) (X ≤ 300) = 20 + 10 + 25 = 55%

Therefore, percentage of workers who makes less than or equal to 300 dollars = 55%

ii) (300 ≤ X ≤ 500) = 30 + 5 = 35%

Therefore, percentage of workers who makes between 300 and 500 dollars = 35%

iii) (400 ≤ X ≤ 700) = 5 + 5 + 5 = 15%

Therefore, percentage of workers who makes between 400 and 700 dollars = 15%